# Differential Privacy of Cross-Attention with Provable Guarantee

### v1

@@
Jiuxiang Gu jigu@adobe.com. Adobe Research.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song zsong@adobe.com. Adobe Research.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
Cross-attention has become a fundamental module nowadays in many important artificial intelligence applications, e.g., retrieval-augmented generation (RAG), system prompt, guided stable diffusion, and many so on. Ensuring cross-attention privacy is crucial and urgently needed because its key and value matrices may contain sensitive information about companies and their users, many of which profit solely from their system prompts or RAG data. In this work, we design a novel differential privacy (DP) data structure to address the privacy security of cross-attention with a theoretical guarantee. In detail, let $n$ be the input token length of system prompt/RAG data, $d$ be the feature dimension, $0<\alpha\leq 1$ be the relative error parameter, $R$ be the maximum value of the query and key matrices, $R_{w}$ be the maximum value of the value matrix, and $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods. Then, our data structure requires $\widetilde{O}(ndr^{2})$ memory consumption with $\widetilde{O}(nr^{2})$ initialization time complexity and $\widetilde{O}(\alpha^{-1}r^{2})$ query time complexity for a single token query. In addition, our data structure can guarantee that the user query is $(\epsilon,\delta)$-DP with $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}R^{2s}R_{w}r^{2})$ additive error and $n^{-1}(\alpha+\epsilon_{s})$ relative error between our output and the true answer. Furthermore, our result is robust to adaptive queries in which users can intentionally attack the cross-attention system. To our knowledge, this is the first work to provide DP for cross-attention. We believe it can inspire more privacy algorithm design in large generative models (LGMs).
@@ 1 Introduction
One fundamental technique used in LGMs is cross-attention [86], which is an essential module in retrieval-augmented generation (RAG) [62], system prompt, guided stable diffusion, and many so on. In RAG, to be more professional, the LGMs can answer user input queries by using a domain-specific database under cross-attention, which may contain specific privacy data and knowledge so that the LGMs gain additional power. For system prompts, based on cross-attention, some customized long prompts, e.g., user information or concrete rules, are concatenated before user input to follow human instructions better, which are commonly used in ChatGPT [32], Claude3 [4] and other commercial LGMs.
@@
To fundamentally preserve cross-attention privacy, we can borrow the powerful tools from differential privacy (DP) [17], which provides measurable privacy and combines with statistical machine learning seamlessly [73]. Thus, in this work, we would like to ask and answer the following question,
@@
Our work demonstrates that the cross-attention computation is equivalent to computing the weighted distance problem.
Definition 1.1 (Cross-attention).
Let $n$ and $m$ be the token length of the system prompt/RAG data and input query, respectively. Let $d$ be the feature dimension. Let $\phi:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be any function that measures vector similarity. Given fixed key matrix $K\in[0,R]^{n\times d}$ and fixed value matrix $V\in[-R_{w},R_{w}]^{n\times d}$, for any input query matrix $Q\in[0,R]^{m\times d}$, the goal of the Cross-Attention Computation is to get the matrix $\mathrm{Attn}(Q,K,V)\in\mathbb{R}^{m\times d}$, which is defined as
@@
where $A\in\mathbb{R}^{m\times n}$ and $D\in\mathbb{R}^{m\times m}$ is a diagonal matrix, i.e.,
$A_{i,j}:=\phi(Q_{i},K_{j}),\quad\text{and}\quad D:=\mathrm{diag}(A{\bf 1}_{n}).$
Our main theorem, presented below, provides a robust solution of cross-attention, ensuring privacy and accuracy guarantees.
@@
Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma 6.7). Let $l=\widetilde{O}(r)$. Let $\phi$ be the softmax distance function (Definition 4.1). Then, our Alg. 4 requires $O(lndr)$ memory to make cross-attention DP with $O(lnr)$ initialization time and $\widetilde{O}(\alpha^{-1}lr)$ query time, such that with probability $1-p_{f}$, the output is $(\epsilon,\delta)$-DP and robust to adaptive query with relative error $n^{-1}(\alpha+\epsilon_{s})$ and additive error $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}lR^{2s}R_{w}r)$.
@@
Our main technique in Theorem 1.2 makes cross-attention differentially private by using the polynomial kernel approximation method and transforming it into a weighted distance problem. Then, the above problem can be solved by summing over weighted distances (depending on the value embedding) between the query embedding and the key embedding. We build a data structure for weighted softmax queries in Section 4.2, and we extend this data structure to handle adaptive queries using the $\epsilon_{0}$-net/metric entropy argument in Section 4.3.
Our contributions can be summarized as follows:
@@
  • We demonstrate that cross-attention computations are equivalent to the weighted distance problem (Section 3.4).
@@
  • We design a novel algorithm (Alg. 3) that can privately answer weighted softmax queries with high probability and a concrete accuracy bound.
@@
  • Our algorithm (Alg. 4) can handle multiple queries and is robust against adaptive query attacks (Theorem 4.4), meaning that potential attackers cannot intentionally extract information of system prompts/RAG data.
@@
To our knowledge, this is the first work to utilize DP to protect prompts in LGMs with theoretically provable guarantees. While some papers have explored protecting user/system prompts with DP [24, 66], they are primarily empirical and lack theoretical guarantees. Additionally, many others are working on protecting private datasets by applying DP to the fine-tuning stage of LGMs training [7, 77, 59, 96, 63, 81], which diverges from our work. The strength of DP lies in its strong, unambiguous, and concrete definition of privacy, enabling algorithm designs with provable privacy and accuracy analysis. Therefore, we believe that the theoretical aspects of DP applications in LGMs remain a highly impactful direction, and we aim to pave the way for further exploration in this area.
Roadmap.
In Section 2, we provide a review of related literature. In Section 3, we present the preliminary of differential privacy (DP) and cross-attention. In Section 4, we outline the main results of our algorithms. In Section 5, we present the main result of our cross-attention theorem (Theorem 5.1). In Section 6, we offer an overview of our techniques. In Section 7, we discuss DP-related topics and potential extensions. In Section 8, we conclude our paper.
2 Related Work
2.1 Differential Privacy Guarantee Analysis
Ever since [17] proposes the notion of differential privacy (DP), it has become one of the most essential standards of privacy protection in both theoretical and empirical ways [20, 61, 98, 73, 94]. DP provides a powerful, robust, and quantifiable privacy definition, allowing algorithm design with concrete privacy and accuracy guarantee [22, 3, 57, 49, 35, 8, 10, 23, 14, 46, 67, 68, 53, 58, 26, 27, 56, 15, 11, 16, 25, 42, 37, 60, 36, 21, 84, 19, 91, 83, 44]. Additionally, new mechanisms have been proposed beyond the traditional Laplace, Gaussian, and Exponential mechanisms [18]. For example, truncated Laplace mechanism [30] is proved to be the current tightest the lower and upper bounds on the minimum noise amplitude and power cross all $(\epsilon,\delta)$-DP distributions.
2.2 Differential Privacy in Data Structure and Attention
Differential privacy (DP) is a flourishing and powerful technique that has enormous applications in the topic of private machine learning. In the era of Large Generative Models (LGMs), there are three primary approaches to ensuring privacy: (1) during the pre-training stage: to protect training data [2, 73], (2) during the adaptation stage: to protect target data [7, 77, 59, 96, 63, 81], (3) during the inference stage: to protect user/system prompts [24] and RAG data [62].
@@ 2.2 Differential Privacy in Data Structure and Attention
To protect training data, DP-SGD [2] uses DP optimizer to ensure data privacy, severing as the traditional baseline method. Recently, numerous works have aimed to improve this method by integrating DP in both the pre-training and fine-tuning stages of LGMs [96, 63, 29, 7, 81, 65, 77, 99, 59]. To protect user/system prompts, [24] provides a survey on both DP and non-DP methods. In the use of LGMs, prompting methods almost become a standard way for inference [80]. Given the billions of prompt interactions daily, ensuring privacy is essential [66].
2.3 Cross-Attention in System Prompt, RAG, Stable Diffusion and More
Cross-attention [86], first introduced in language translation, is a widely used technique in many advanced AI systems. For example, Stable Diffusion [76] and SORA [70] employ cross-attention as a core module for a text-to-image conditional generation. This technique is also utilized by other multimodal models [40], including Imagen [79] and Diffusion Transformer [74].
In the realm of text-to-image editing, [48] analyzes and controls the cross-attention module to enable editing without requiring additional training. Furthermore, [97] tackles the issue of inaccurate cross-attention maps, enhancing fine-grained control over edited regions while preventing unintended changes to other areas.
In addition, Retrieval Augmented Generation (RAG) [62, 9, 45], a technique that improves model responses by retrieving information from a knowledge base or external documents, extensively uses cross-attention as its core design module.
Cross-attention also has other applications. [71] demonstrates that the prompt-tuning [39] task can be formulated as cross-attention, while [12] uses cross-attention to fuse multi-scale features in vision transformers, thereby reducing computation. Moreover, attention-based Transformer architecture makes LGMs equipping many emergent ability [89], such as spatial reasoning [87], mathematical reasoning [38], in-context learning ability [82], compositional ability [92], few-shot adaptation ability [78, 93], and so on.
@@ 3 Preliminary
In this section, we give the preliminary of differential privacy (DP) and cross-attention. In Section 3.1, we describe the notations we use. In Section 3.2, we give several definitions related to DP. In Section 3.3, we give the DP facts we use in the paper. In Section 3.4, we show that the cross-attention module is equivalent to the weighted distance problem.
@@ 3.2 Differential Privacy Definitions
In this section, we give several definitions related to differential privacy (DP). We refer the reader to [18] for more background and details on DP.
@@
The sensitivity of a function $f:\mathbb{R}^{n}\to\mathbb{R}^{d}$ is defined by:
$\Delta:=\max_{X,X^{\prime}\in\mathbb{R}^{n},\|X-X^{\prime}\|_{1}=1}\|f(X)-f(X^{\prime})\|_{1}.$
@@
For the Laplace mechanism, we have the following definitions.
Definition 3.4 (Laplace distribution).
We use $\mathrm{Lap}(b)$ to denote the pdf:
$p(z)=\frac{1}{2b}\exp(-\frac{|z|}{b}).$
Fact 3.5.
@@
Lemma 3.8 (Laplace mechanism, [18, 30], see Lemma 2.2 in [3]).
@@
3.3 DP Facts
Fact 3.9 (Post-processing, see Fact 2.1 in [35]).
Fact 3.10 (Basic composition, see Fact 2.3 in [35]).
We can do much better if we know that the inputs are disjoint.
Fact 3.11 (Parallel composition, see Fact 2.4 in [35]).
Theorem 3.12 (Advanced composition, see Theorem 3.20 in [18]).
Comparison of truncated Laplace, Gaussian, and Laplace mechanisms
In this paper, we use the Chebyshev inequality to bound the error, and from [30], we know that the truncated Laplace mechanism has the current minimum variance across all $(\epsilon,\delta)$-DP distributions.
The variance of Gaussian mechanism in Theorem 3.22 in [18]:
The variance of Laplace mechanism in Fact 3.5:
$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}}{\epsilon^{2}}.$
The variance of truncated Laplace mechanism in Fact 3.7, for $c\in(0,1]$:
$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}c}{\epsilon^{2}}.$
3.4 Cross-Attention
Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1. We can access $K$ and $V$ before cross-attention inference but not user input $Q$. For the attention mechanism $\mathrm{Attn}$ (Definition 1.1), matrix $AV$ can be made DP.
Now, we describe how we can use our techniques to solve this problem. Let $A_{i,j}=\phi(Q_{i},K_{j})$ for $i\in[m],j\in[n]$. Let $V_{j,k}\in\mathbb{R}$ be the $(j,k)$-th entry of $V$, for $j\in[n],k\in[d]$.
In the following discussion, we focus on $\phi$ being the exponential inner product/softmax case, so we have $\phi(Q_{i},K_{j})=\exp(\langle Q_{i},K_{j}\rangle/d)$.
By post-processing (Fact 3.9), we only need to make $AV$ DP, where the $(i,k)$-th entry of $AV$ for each $i\in[m],k\in[d]$ can be computed by
$(AV)_{i,k}=$ $~{}\sum_{j=1}^{n}\phi(Q_{i},K_{j})V_{j,k}=~{}\sum_{j=1}^{n}V_{j,k}\exp(\langle Q_{i},K_{j}\rangle/d),$
which can be viewed as a weighted softmax problem, where $V$ is the weights, $Q$ is the query, and $K$ is the dataset. Thus, we only need to add noise to $K$. Furthermore, we find that we can only handle one column of $V$, $V_{*,k}\in\mathbb{R}^{n}$, in one data structure. Therefore, we need to initialize a total of $d$ different DPTreeSoftmaxAdaptive (Alg. 4) structures, each with weights $V_{*,k}$ for $k\in[d]$.
Algorithm 1 DPTree initialization function
1:datastructure DPTree $\triangleright$ Theorem 6.1
2:members
3:     $a:\mathbb{R}^{n}$
4:     $b:\mathbb{R}^{2n-1}$
5:     $c:\mathbb{R}^{2n-1}$
6:end members
7:procedure Init($a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1)$) $\triangleright$ Lemma B.3, Lemma 6.3
8:     $a\leftarrow a$
9:     $b[n,2n-1]\leftarrow a$
10:     for $i=n\to 2n-1$ do
11:         $c[i]\leftarrow b[i]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
12:     end for
13:     for $i=(\log n)\to 1$ do
14:         for $j=1\to 2^{i-1}$ do
15:              $k\leftarrow 2^{i-1}+j-1$
16:              $b[k]\leftarrow b[2k]+b[2k+1]$
17:              $c[k]\leftarrow b[k]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
18:         end for
19:     end for
20:end procedure
21:end datastructure
4 Main Results: DPTree
In this section, we provide the main results for our data structures: DPTree (Algorithm 1 and 2), DPTreeSoftmax (Algorithm 3), and DPTreeSoftmaxAdaptive (Algorithm 4). In Section 4.1, we give our basic building block algorithm DPTree. In Section 4.2, we present ourDPTreeSoftmax algorithm that can solve the weighted softmax problem. In Section 4.3, we present our DPTreeSoftmaxAdaptive algorithm that enables DPTreeSoftmax to handle adaptive query problem.
Algorithm 2 DPTree query
1:datastructure DPTree $\triangleright$ Theorem 6.1
2:procedure Query($x\in[n],y\in[n]$) $\triangleright$ Lemma B.4, B.5, B.6
3:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $c$) by using at most $2\log n$ nodes on the path. Let Value be the above summation.
4:     return Value
5:end procedure
6:procedure TrueQuery($x\in[n],y\in[n]$)
7:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $b$) by using at most $2\log n$ nodes on the path. Let Value be the above summation.
8:     return Value
9:end procedure
10:end datastructure
4.1 DPTree Initialization and Query
We design a basic data structure DPTree (Algorithm 1 and 2) that can answer summation queries by a summation segment tree with truncated Laplace noise. The algorithm first builds a binary summation tree in an array and then adds truncated Laplace noises to each node (Algorithm 1).
In query time, we first trace from bottom nodes to find their lowest common ancestor, then report the summation by using at most $O(\log n)$ nodes on the path (Algorithm 2).
Algorithm 3 Softmax query
1:datastrucutre DPTreeSoftmax $\triangleright$ Theorem 4.2
2:members
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{r}:\textsc{DPTreeDistance}$triangleright Alg. 6, Theorem G.7
4:     $P:[0,\Gamma_{R,s}]^{n\times r}$triangleright Def. G.3 for $\Gamma_{R,s}$, Eq. (6) for $s$, Eq. (7) for $r$, and Lem. G.5 for $P$
5:     $w:[-R_{w},R_{w}]^{n}$
6:     $P_{wx},~{}s_{w},~{}\epsilon_{s}:\mathbb{R}$
7:end members
8:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$, $\epsilon_{s}\in(0,0.1)$) $\triangleright$ Lemma 6.7
10:     for $j=1\to n$ do
11:         Compute $P(x_{j})$
12:         Compute $w_{j}\|P(x_{j})\|_{2}^{2}$
13:         $P_{wx}\leftarrow P_{wx}+w_{j}\|P(x_{j})\|_{2}^{2}$
14:         $s_{w}\leftarrow s_{w}+w_{j}$
15:         $P_{j,:}\leftarrow P(x_{j})$
16:     end for
17:     for $i=1\to r$ do
18:         $\mathcal{D}_{i}.$Init($P_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{r\log(1/\delta^{\prime})}$, $\delta/r$) $\triangleright$ Alg. 6
19:     end for
20:end procedure
21:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma 6.7
22:     Value $\leftarrow 0$
23:     Compute $P(y)$
24:     Compute $\|P(y)\|_{2}^{2}$
25:     for $i=1\to r$ do
26:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(P(y)_{i},\alpha)$triangleright Alg. 7
27:     end for
29:     return Value
30:end procedure
31:end datastrucutre
@@ 4.1 DPTree Initialization and Query
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(r\log(dR/(\epsilon_{s}p_{f})))}:\textsc{DPTreeSoftmax}$triangleright Alg. 3

### v2

@@
Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
Cross-attention has become a fundamental module nowadays in many important artificial intelligence applications, e.g., retrieval-augmented generation (RAG), system prompt, guided stable diffusion, and many more. Ensuring cross-attention privacy is crucial and urgently needed because its key and value matrices may contain sensitive information about model providers and their users. In this work, we design a novel differential privacy (DP) data structure to address the privacy security of cross-attention with a theoretical guarantee. In detail, let $n$ be the input token length of system prompt/RAG data, $d$ be the feature dimension, $0<\alpha\leq 1$ be the relative error parameter, $R$ be the maximum value of the query and key matrices, $R_{w}$ be the maximum value of the value matrix, and $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods. Then, our data structure requires $\widetilde{O}(ndr^{2})$ memory consumption with $\widetilde{O}(nr^{2})$ initialization time complexity and $\widetilde{O}(\alpha^{-1}r^{2})$ query time complexity for a single token query. In addition, our data structure can guarantee that the process of answering user query satisfies $(\epsilon,\delta)$-DP with $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}R^{2s}R_{w}r^{2})$ additive error and $n^{-1}(\alpha+\epsilon_{s})$ relative error between our output and the true answer. Furthermore, our result is robust to adaptive queries in which users can intentionally attack the cross-attention system. To our knowledge, this is the first work to provide DP for cross-attention and is promising to inspire more privacy algorithm design in large generative models (LGMs).
@@ 1 Introduction
One fundamental technique used in LGMs is cross-attention [97], which is an essential module in retrieval-augmented generation (RAG) [68], system prompt, guided stable diffusion, and many so on. In RAG, to be more professional, the LGMs answer user input queries by using a domain-specific database under cross-attention, which may contain specific privacy data and knowledge so that the LGMs gain additional power. For system prompts, based on cross-attention, some customized long prompts, e.g., user information or concrete rules, are concatenated before user input to follow human instructions better, which are commonly used in ChatGPT [34], Claude3 [4] and other commercial LGMs.
@@
To fundamentally preserve cross-attention privacy, we borrow the powerful tools from differential privacy (DP) [19], which provides measurable privacy and combines with statistical machine learning seamlessly [84]. Thus, in this work, we would like to ask and answer the following question,
@@
Our work demonstrates that the Softmax cross-attention computation is equivalent to computing the weighted distance problem.
Definition 1.1 (Softmax cross-attention).
Let $n$ and $m$ be the token length of the data and input query, respectively. Let $d$ be the feature dimension. Given fixed key matrix $K\in[0,R]^{n\times d}$ and fixed value matrix $V\in[-R_{w},R_{w}]^{n\times d}$, for any input query matrix $Q\in[0,R]^{m\times d}$, the goal of the Softmax Cross-Attention Computation is to get the matrix $\mathrm{Attn}(Q,K,V)\in\mathbb{R}^{m\times d}$, which is
@@
where $A\in\mathbb{R}^{m\times n}$ satisfies $A_{i,j}:=\exp(\langle Q_{i},K_{j}\rangle/d)$ for any $i\in[m],j\in[n]$ ($Q_{i}$ and $K_{j}$ denote the $i$-th and $j$-th rows of $Q$ and $K$, respectively) and $D:=\mathrm{diag}(A{\bf 1}_{n})\in\mathbb{R}^{m\times m}$ is a diagonal matrix.
Note that $\mathsf{Softmax}(QK^{\top})=D^{-1}A\in\mathbb{R}^{m\times n}$ in Definition 1.1, which is the standard function used in transformers, and usually, we call it as attention matrix. Our main theorem, presented below, provides a robust solution of cross-attention, ensuring privacy and accuracy guarantees.
@@
Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma C.7). Let $l=\widetilde{O}(r)$. Then, our Algorithm 1 requires $O(lndr)$ memory with $O(lnr)$ initialization time and $\widetilde{O}(\alpha^{-1}lr)$ query time, such that with probability $1-p_{f}$, the output process of cross-attention satisfies $(\epsilon,\delta)$-DP and is robust to adaptive query with relative error $n^{-1}(\alpha+\epsilon_{s})$ and additive error $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}lR^{2s}R_{w}r)$.
@@
Our main technique in Theorem 1.2 ensures that cross-attention is differentially private by using the polynomial kernel approximation method and transforming it into a weighted distance problem. We then solve the problem by summing over weighted distances (depending on the value embedding) between the query embedding and the key embedding. We build a data structure for weighted Softmax queries in Section 4.3, and we extend this data structure to handle adaptive queries using the $\epsilon_{0}$-net/metric entropy argument in Section 4.4. Furthermore, our error decreases as the input token length grows, diminishing both the relative and additive errors to zero.
Our contributions are as follows:
@@
  • We demonstrate that cross-attention computations are equivalent to the weighted distance problem (Section 3).
@@
  • We design a novel algorithm (Algorithm 3) that privately answers weighted Softmax queries with high probability and a concrete accuracy bound.
@@
  • Our algorithm (Algorithm 1) handles multiple cross-attention queries and is robust against adaptive query attacks (Theorem 3.1), meaning that potential attackers cannot intentionally extract information of system prompts/RAG data.
@@
To our knowledge, this is the first work to utilize DP to protect prompts in LGMs with theoretically provable guarantees. While some have explored protecting user/system prompts with DP [26, 77], they are primarily empirical and lack theoretical guarantees. Additionally, many others are working on protecting private datasets by applying DP to the fine-tuning stage of LGMs [9, 88, 63, 113, 74, 92], which diverges from our work. The strength of DP lies in its strong, unambiguous, and concrete definition of privacy, enabling algorithm designs with provable privacy and accuracy analysis. Therefore, we believe that the theoretical aspects of DP applications in LGMs remain a highly impactful direction, and we aim to pave the way for further exploration in this area.
1.1 Related Work
@@
Differential Privacy in Data Structure and Attention. Differential privacy (DP) is a flourishing and powerful technique that has enormous applications in the topic of private machine learning. In the era of Large Generative Models (LGMs), there are three primary approaches to ensuring privacy: (1) during the pre-training stage: to protect training data [2, 84], (2) during the adaptation stage: to protect target data [9, 88, 63, 113, 74, 92], (3) during the inference stage: to protect user/system prompts [26] and RAG data [68]. To protect training data, DP-SGD [2] uses DP optimizer to ensure data privacy, severing as the traditional baseline method. Recently, numerous works have aimed to improve this method by integrating DP in both the pre-training and fine-tuning stages of LGMs [113, 74, 31, 9, 92, 76, 88, 116, 63]. To protect user/system prompts, EW [26] provides a survey on both DP and non-DP methods. In the use of LGMs, prompting methods almost become a standard way for inference [91]. Given the billions of prompt interactions daily, ensuring privacy is essential [77]. We refer readers to Appendix A for more related works.
Roadmap. In Section 2, we present the preliminary of differential privacy (DP) and cross-attention. In Section 3, we present the main result of our cross-attention theorem (Theorem 3.1). In Section 4, we outline the main results of our algorithms. In Section 5, we discuss DP-related topics and potential extensions. In Section 6, we conclude our paper.
@@ 2 Preliminary
In this section, we give the preliminary of differential privacy (DP) and cross-attention. In Section 2.1, we describe the notations. In Section 2.2, we give definitions related to DP.
@@ 2.2 Differential Privacy Definitions
In this section, we give several definitions related to differential privacy (DP). We refer the reader to DR [20] for more background and details on DP.
@@
The sensitivity of a function $f:\mathbb{R}^{n}\to\mathbb{R}^{d}$ is defined by: $\Delta:=\max_{X,X^{\prime}\in\mathbb{R}^{n},\|X-X^{\prime}\|_{1}=1}\|f(X)-f(X^{\prime})\|_{1}.$
@@
We mainly use the truncated Laplace mechanism, which has the following definitions.
@@
Lemma 2.6 (Laplace mechanism, [20, 32], see Lemma 2.2 in AIMN [3]).
@@
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(r\log(dR/(\epsilon_{s}p_{f})))}:\textsc{DPTreeSoftmax}$triangleright Algorithm 3